# arXiv:2603.02085v1[hep-ph] 02 Mar 2026

### v1

@@ VI Summary
In summary, to investigate the limiting temperature of the hadron resonance gas model with EVE, we have studied the relation between the effective degrees of freedom and the trace anomaly in the HRG model. If we regard the thermodynamic relation as the evolution equation and define the effective degree of freedom as $P/T^{4}$ (or $P/\mu^{4}$), we obtain the equations that relate to the trace anomaly. The structures of the equations resemble the equation of the so called c-theorem in the two dimensional conformal field theory. There is a stationary point where the (modified) trace anomaly vanishes and the scale invariance is restored.
We consider two types of the c-theorem like condition for the effective degrees of freedom. The first condition is that the EDOF should not decrease as the energy scale increases. If we use the $(T,\mu)$ coordinate, according to this condition, the modified trace anomaly should not be negative. However, the limiting temperature obtained by this requirement is much higher than the crossover transition temperature predicted by the LQCD calculation  Borsanyi et al. (2020), and, at $\mu=0$, it is close to the transition temperature $T_{\rm d}\sim 0.285$ GeV in the pure gluonic theory Borsanyi et al. (2022); Giusti et al. (2025); Fujimoto et al. (2025). If we use the $(\beta,\gamma)$ coordinate, $T_{\rm zero}^{\prime}$ where $\Delta_{\rm B}$ vanishes at $\gamma=0$ is also close to $T_{\rm d}$.
@@
If we use the $(T,\mu)$ coordinate, according to this condition, the $T^{5}$-normalized trace anomaly should not decrease as $T$ increases. The temperature at which the modified trace anomaly reaches its maximum is the limiting temperature of the HRG model with EVE. The obtained limiting temperature of the HRG model with EVE (EVE2) is consistent with the one obtained by using the normalized baryon fluctuation in the previous work and the LQCD predicted critical point Shah et al. (2024).
@@
If we use the $(\beta,\gamma)$ coordinate, $T_{\rm max}^{\prime}$ where $\beta^{3}\Delta_{\rm B}$ has its maximum for the fixed value of $\gamma$ is somewhat higher than $T_{\rm max}$ obtained by using the $(T,\mu)$ coordinate but is close to the curve determined by the condition $n_{\rm b}=1/(2v_{\rm B})$. In particular, at $\mu=0$, $T_{\rm max}^{\prime}$ almost coincides with the Roberge-Weiss like temperature $T_{\rm RWL}$ Oshima et al. (2026) which is the limiting temperature of the HRG model with EVE at $\mu=i(2k+1)\pi T$ where $k$ is any integer. This indicates that the condition (15) is weaker than (13).
@@
Of course, these limiting temperatures are for the HRG with EVE but not for QCD. In fact, at $\mu=0$, the EDOF is convex upwards in the high temperature region. Hence, hadron matter is expected to transition to quark matter or exotic new matter in the high temperature region.
@@
There is an open question as to why the temperature of the maximum of the $T^{5}$-scaled modified trace anomaly is so close to that of the normalized baryon number fluctuation. It is important that these temperatures are just below the temperature determined by the condition $n_{\rm b}=1/(2v_{\rm B})$. This indicates that the suppression factor $1/(1+v_{\rm B}n_{\rm b})$ plays a crucial role. Detailed study of the role of the suppression factor is needed in the future.
@@
It is also very interesting that, in the HRG model with EVE2, the curve of the maximum $\Delta-\mu n_{\rm B}$ crosses with the curve of the curve of $\chi_{2}^{\rm B}$, where the normalized baryon fluctuation has its maximum, near the CP predicted by the LQCD calculation Shah et al. (2024). This indicates that the limitation of the HRG model with EVE has some information about the critical point, although the model has no mechanism for the chiral phase transition. It was pointed out that, near the critical point, the ordering density is a linear combination of the scalar density, the net baryon number density, and the energy density rather than the pure scalar density itself Fujii and Ohtani (2004). The effects of the baryon number density and the trace anomaly may be more important in the large $\mu$ region. Further study to investigate the nature of the transition beyond CP is needed.

### v2

@@ VI Summary
In summary, to investigate the limiting temperature of the hadron resonance gas model with EVE, we have studied the relation between the effective degrees of freedom and the trace anomaly in the HRG model. If we regard the thermodynamic relation as the evolution equation and define the effective degree of freedom as $P/T^{4}$ (or $P/\mu^{4}$), we obtain the equations that relate to the trace anomaly. The structures of the equations resemble the equation of the so-called c-theorem in the two dimensional conformal field theory. There is a stationary point where the (modified) trace anomaly vanishes and the scale invariance is restored.
We consider two types of the c-theorem like condition for the effective degrees of freedom of the HRG model with EVE. The first condition requires that the EDOF should not decrease as the energy scale increases. If we use the $(T,\mu)$ coordinate, according to this condition, the modified trace anomaly should not be negative. However, the limiting temperature obtained by this requirement is much higher than the crossover transition temperature predicted by the LQCD calculation  Borsanyi et al. (2020), and, at $\mu=0$, it is close to the transition temperature $T_{\rm d}\sim 0.285$ GeV in the pure gluonic theory Borsanyi et al. (2022); Giusti et al. (2025); Fujimoto et al. (2025). If we use the $(\beta,\gamma)$ coordinate, $T_{\rm zero}^{\prime}$ where $\Delta_{\rm B}$ vanishes at $\gamma=0$ is also close to $T_{\rm d}$.
@@
If we use the $(T,\mu)$ coordinate, according to this condition, the $T^{5}$-normalized trace anomaly should not decrease as $T$ increases. The temperature at which the modified trace anomaly reaches its maximum is the limiting temperature of the HRG model with EVE. The obtained limiting temperature of the HRG model with EVE (EVE2) is consistent with the one obtained by using the normalized baryon fluctuation in the previous study and the LQCD predicted critical point Shah et al. (2024). This result supports our assumption that the temperature of the maximum $\chi_{2}^{\rm B}$ is the limiting temperature of the HRG model with EVE.
@@
If we use the $(\beta,\gamma)$ coordinate, $T_{\rm max}^{\prime}$ where $\beta^{3}\Delta_{\rm B}$ has its maximum for the fixed value of $\gamma$ is somewhat higher than $T_{\rm max}$ obtained by using the $(T,\mu)$ coordinate but is close to the curve determined by the condition $n_{\rm b}=1/(2v_{\rm B})$. In particular, at $\mu=0$, $T_{\rm max}^{\prime}$ almost coincides with the Roberge-Weiss like temperature $T_{\rm RWL}$ Oshima et al. (2026) which is the limiting temperature of the HRG model with EVE at $\mu=i(2k+1)\pi T$ where $k$ is any integer. This indicates that the condition (15) is weaker than (13). There is an open question which coordinate system is most appropriate to determine the limiting temperature. A more fundamental and microscopic understanding of the increase and decrease in degrees of freedom is needed.
@@
It should be also remarked that these limiting temperatures are for the HRG with EVE but not for QCD. In fact, at $\mu=0$, the EDOF is convex upwards in the high temperature region. Hence, hadron matter is expected to transition to quark matter or exotic new matter in the high temperature region. Furthermore, since our transition scenario is based on the observation on the crossover transition at $\mu=0$, it is an open question whether these limiting conditions are satisfied or not, beyond the critical point where the fluctuations shows the divergent behavior.
@@
There is also an open question why the temperature of the maximum of the $T^{5}$-scaled modified trace anomaly is so close to that of the normalized baryon number fluctuation. It is important that these temperatures are just below the temperature determined by the condition $n_{\rm b}=1/(2v_{\rm B})$. This indicates that the suppression factor $1/(1+v_{\rm B}n_{\rm b})$ plays a crucial role. The condition $n_{\rm b}=1/(2v_{\rm B})$ is also related to the $R=0$ criterion in the thermodynamic geometry Oshima et al. (2026). Detailed study of the role of the suppression factor is needed in the future.
@@
It is very interesting that, in the HRG model with EVE2, the curve of the maximum $\Delta-\mu n_{\rm B}$ crosses with the curve of the maximum $\chi_{2}^{\rm B}$ near the CP predicted by the LQCD calculation Shah et al. (2024). This indicates that the limitation of the HRG model with EVE has some information about the critical point, although the model has no mechanism for the chiral phase transition. It was pointed out that, near the critical point, the ordering density is a linear combination of the scalar density, the net baryon number density, and the energy density rather than the pure scalar density itself Fujii and Ohtani (2004). The effects of the baryon number density and the trace anomaly may be more important in the large $\mu$ region. Further study to investigate the nature of the transition beyond CP is needed.